OpenStudy (anonymous):

17-12=1/2(a)+3/4(a)-5/8(a)

OpenStudy (anonymous):

-5= 4/8(a) + 6/8(a) - 5/8(a), -5=10/8(a)-5/8(a), -5= 5/8(a), a= -8

OpenStudy (anonymous):

o, it's 8 nvm

OpenStudy (anonymous):

set out common denominators and then add all the a's together and then set equal to 17-12=5

OpenStudy (anonymous):

i got 1/8 but its wrong

OpenStudy (anonymous):

it's just positive 8

OpenStudy (anonymous):

exactly i just solved the problem and i got 8=a

OpenStudy (anonymous):

yea, hi-5 hahaha

OpenStudy (anonymous):

lol =]